noahcam1349:

Kevin needs to run errands today, including a trip to the gas station, the grocery store, and the dry cleaner. If he begins at the gas station, continues to the dry cleaner, then to the grocery store, and ends up back at the gas station again, how many miles will he have traveled in total?

Spectrum:

Well, that is undeetectabe.

noahcam1349:

In order to find out how many miles in total Kevin has traveled, you will need to add up all the sides of the triangle. Dry cleaner to gas station: 4 miles Gas station to grocery store: 8 miles However, to find the distance from the dry cleaner to the grocery store, you will need to use the Pythagorean Theorem. a2+b2 = c2 42+82 = c2 16+64 = c2 80 = c2 √80 = √c2 8.94 ≈ c Now that you know all the distances, add up the sides of the triangle to see how far Kevin traveled. 4 + 8.94 + 8 = 20.94 miles in all

noahcam1349:

thats the answer

juju2021:

yoooo

noahcam1349:

yo

Olivia144:

@juju2021 heyy